Shift report 20000805 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Aug 05 2000 - 23:58:40 EDT

  • Next message: http account: "Shift report 20000806 00:00-08:00"

    ------------------------------------------------------------
    Date: Sat Aug  5 23:58:39 2000, Shift: 16:00-24:00
    Supervisor: YB, with: TT
    Log book#4 pages 6 - 7
    
    Shift Summary:
    ------------------------------------------------------------
    1. Cogged beam at gamma = 70 for about 2h since 21:50. 
    2. - Runs 2061 and 2062 is still running. Very low rate. 
         Almost no collision events (less than 200 in run 2061).
       - Changed MTPC2 HV from 1230V to 1250V. Saw a few nice 
         tracks.
       - TOFW CAMAC readout is fixed by KO.
    3. - FFS1440 is not responding and were found off. Turned on
         H2 and TOFW.
       - H1 and C1 HV mainframes are not responding. Seem to be 
         off (by SuperMon pictures)
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sun Aug 06 2000 - 00:01:54 EDT